undeadknight26 (undeadknight26):

Hello!!! Part B: Write any two solutions for f(x). (3 points) Graph provided easy medal/ or fan!

undeadknight26 (undeadknight26):

I think it is (4,0) and (5,0)

undeadknight26 (undeadknight26):

@SolomonZelman @shamil98

OpenStudy (solomonzelman):

Solution is a point of intersection, but I don't quite get the question (as always)

OpenStudy (solomonzelman):

I think theya e asking for the equation.

undeadknight26 (undeadknight26):

me neither...maybe...@phi lets see what this guy says...

OpenStudy (solomonzelman):

\[f(x)=2x-5\]

OpenStudy (solomonzelman):

I found the slope, and the y-intercept is shown.

OpenStudy (shamil98):

yeah, solomon has it right

OpenStudy (shamil98):

a solution for f(x) could be any point on that line

OpenStudy (shamil98):

(1,-3) for example

undeadknight26 (undeadknight26):

so not for this i plug in a a number and solve?

OpenStudy (solomonzelman):

what do you mean?

OpenStudy (shamil98):

it's asking for any two solutions right?

OpenStudy (shamil98):

you have f(x) = 2x - 5 solve for when x = 0 , and x = 1 to get two solutions

undeadknight26 (undeadknight26):

oh ok...i understand so: 4, 3 2,-1?

OpenStudy (solomonzelman):

YES

OpenStudy (shamil98):

f(0) = 2(0) -5 = -5 f(1) = 2(1)- 5 f (1) = -3 so two solutions could be (0,5) , and (1,-3)

OpenStudy (shamil98):

yes those are also correct

undeadknight26 (undeadknight26):

can i use the ones i mentioned?

OpenStudy (solomonzelman):

uk, (0, -5)

OpenStudy (shamil98):

oops yeah (0,-5)*

undeadknight26 (undeadknight26):

Now the next and last part says: What is the solution to the equation p(x) = g(x)? Justify your answer. (4 points)

OpenStudy (shamil98):

I would say (1,3) because that's where the two equations intersect.

OpenStudy (solomonzelman):

iYeah, i also thought that.

undeadknight26 (undeadknight26):

i had to tell them the solution for the first part thats why i didn't use them...can you guys help me with the last part? @shamil98 @SolomonZelman

undeadknight26 (undeadknight26):

@phi @PixieDust1

undeadknight26 (undeadknight26):

@agent0smith @ranga @nincompoop

undeadknight26 (undeadknight26):

What is the solution to the equation p(x) = g(x)? Justify your answer.

OpenStudy (phi):

the solution to p(x)= g(x) is the intersection point.

undeadknight26 (undeadknight26):

oh i'sa idiot...lol so it would b 1,3 because etc?

OpenStudy (phi):

yes

undeadknight26 (undeadknight26):

thank you my friend!